# Distinguishing the origin of eccentric black-hole mergers with gravitational-wave spin measurements

### v1

@@ Acknowledgements
We thank the Sexten Center for Astrophysics, where this work was kickstarted. We thank Johan Samsing and Ilya Mandel for discussions. J.S. acknowledges funding from NWO Vidi Grant No. 639.042.728–BinWaves. D.G. and G.F. are supported by ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ). and the ICSC National Research Centre funded by NextGenerationEU. D.G. is supported by MSCA Fellowship No. 101064542–StochRewind, MSCA Fellowship No. 101149270–ProtoBH, and MUR Young Researchers Grant No. SOE2024-0000125. I.M.R.-S. acknowledges the support of the Herchel Smith fund.
@@ Appendix A Evolution of $\chi_{\rm eff}$ during eccentric BBH inspirals
During the inspiral of a BBH, the effective spin parameter $\chi_{\rm eff}$ is a conserved quantity at second post-Newtonian (PN) order when averaging over orbital motion (Racine, 2008; Gerosa et al., 2023). On highly eccentric orbits the orbit-averaged approximation may break down if efficient GW emission at close periapsis passages induces changes of the orbital elements on a timescale that subceeds the orbital period. As a result, the black-hole spins may precess about the orbital angular momentum vector without conserving $\chi_{\rm eff}$, so that our method to use $\chi_{\rm eff}$ at BBH formation as a proxy for $\chi_{\rm eff}$ at BBH merger would break down. In order to estimate our error, we simulate the inspiral of one thousand BBHs using the publicly available regularised direct $N$-body integrator MSTAR (Rantala et al., 2020; Mannerkoski et al., 2023) which uses the Gragg-Bulirsch-Stoer extrapolation technique (Gragg, 1965; Bulirsch & Stoer, 1966) to account for the mutual gravitational interaction between the $N=2$ black holes to 3.5PN order (Mora & Will, 2004). The integration is terminated if the BBHs get as close as twelve combined Schwarzschild radii. The BBHs are initialised on a highly eccentric orbit with eccentricity $e=0.999$, total mass $m=30\,\rm M_{\odot}$, and periapsis $r_{p}=300M$ where $M=Gm/c^{2}$. The mass ratio of the BBHs is drawn from a power-law distribution $p(q)\sim q^{\alpha}$ between 0.1 and 1.0 and the spin magnitudes from some Beta distribution $p(\chi_{1,2})\sim\rm Beta(a,b)$ with our fiducial parameters $\alpha=1.1$, $a=1.4$, and $b=3.6$. The BBHs are initialised at apoapsis and their spins are assumed to be aligned. Figure 6 shows that for the majority of BBH inspirals $\chi_{\rm eff}$ does not oscillate by more than $\sim 0.02$ and none of them by more than $\sim 0.1$ around the initial values. Based on these simulations, we conclude that our results are not significantly affected by treating $\chi_{\rm eff}$ as a conserved quantity.
@@ Appendix B Evolution of $\cos\theta_{1}$ and $\cos\theta_{2}$ during eccentric BBH inspirals
Unlike the effective spin parameter $\chi_{\rm eff}$, the spin tilt angles do not remain constant during the inspiral. Instead, they precess within precession cones whose amplitudes evolve due to GW emission, while remaining effectively fixed on the shorter precession timescale. The presence of distinct timescales in the dynamics allows for an efficient treatment of spin evolution through a double-averaging procedure: first over the orbital timescale and then over the precession timescale (Gerosa et al., 2015, 2023). This technique has been successfully applied to study spin-precessing, quasi-circular BBHs and, more recently, eccentric systems as well (Fumagalli & Gerosa, 2023). Using the extended version of the precession code presented by Fumagalli & Gerosa (2023), we investigate the evolution of the spin tilts in a population of eccentric BBHs in the mutually aligned and uncorrelated models as described in Sec. 1. We simulate $10^{5}$ binaries with total mass $M=20M_{\odot}$; the mass ratios are drawn from a power-law distribution $p(q)\sim q^{\alpha}$ with $\alpha=1.1$ and $q\in[0.1,1.0]$; spin magnitudes are either sampled from a Beta distribution $p(\chi_{1,2})\sim\text{Beta}(a,b)$ with $a=1.4$ and $b=3.6$ or fixed to the maximal value $\chi_{1}=\chi_{2}=1$. BBHs are assumed to form with the initial eccentricity $e=0.999$ and semi-major axis $a\sim 1.1\times 10^{5}GM/c^{2}$, corresponding a GW frequency $f_{\rm GW}\sim 0.95$ Hz. We evolve the systems until they reach an eccentricity of $e=0.1$ at $f_{\rm GW}=10$ Hz.
Figure 7: Distribution of the cosine of spin tilt angles, $\cos\theta_{1}$ (left panels) and $\cos\theta_{2}$ (right panels). We assume spin tilts to be either mutually aligned (teal) or uncorrelated (maroon) at formation. Distributions at detection ($f_{\rm GW}=10$ Hz) and BBH formation ($f_{\rm GW}=0.95$ Hz) are shown as filled and dashed histogram, respectively. We assume either maximally spinning BH (top panels) or spin magnitudes extracted from a Beta distribution as in Fig. 2 (bottom panels).
Figure 7 shows the resulting evolution of the spin tilts $\theta_{1,2}$. As previously reported in quasi-circular binaries (Bogdanović et al., 2007; Gerosa et al., 2015), initially isotropic (i.e., uncorrelated) spin tilt distributions remain isotropic throughout the inspiral, even in the presence of eccentricity (Fumagalli & Gerosa, 2023). On the other hand, deviations appear in the case of initially aligned spin tilts. In these systems, the spin tilt distributions evolve away from mutual alignment, exhibiting a tendency to cluster around $\cos\theta_{i}=\pm 1$. This behavior, already identified in the quasi-circular case (Gerosa et al., 2013), is here found to persist in the eccentric regime as well. The effect is especially pronounced in the maximally spinning case, consistent with the expectation that spin-orbit and spin-spin couplings are stronger for higher spin magnitudes.

### v2

@@ Acknowledgements
We thank the Sexten Center for Astrophysics, where this work was kickstarted. We thank Johan Samsing, Ilya Mandel, Fabio Antonini, and the anonymous referee for useful input and discussions. J.S. acknowledges funding from NWO Vidi Grant No. 639.042.728–BinWaves. D.G. and G.F. are supported by ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ). and the ICSC National Research Centre funded by NextGenerationEU. D.G. is supported by MSCA Fellowship No. 101064542–StochRewind, MSCA Fellowship No. 101149270–ProtoBH, and MUR Young Researchers Grant No. SOE2024-0000125. I.M.R.-S. acknowledges the support of the Herchel Smith fund.
@@ Appendix A Evolution of $\chi_{\rm eff}$ during eccentric BBH inspirals
During the inspiral of a BBH, the effective spin parameter $\chi_{\rm eff}$ is a conserved quantity at second post-Newtonian (PN) order when averaging over orbital motion (Racine, 2008; Gerosa et al., 2023). On highly eccentric orbits the orbit-averaged approximation may break down if efficient GW emission at close periapsis passages induces changes of the orbital elements on a timescale that subceeds the orbital period. As a result, the black-hole spins may precess about the orbital angular momentum vector without conserving $\chi_{\rm eff}$, so that our method to use $\chi_{\rm eff}$ at BBH formation as a proxy for $\chi_{\rm eff}$ at BBH merger would break down. In order to estimate our error, we simulate the inspiral of one thousand BBHs using the publicly available regularised direct $N$-body integrator MSTAR (Rantala et al., 2020; Mannerkoski et al., 2023) which uses the Gragg-Bulirsch-Stoer extrapolation technique (Gragg, 1965; Bulirsch and Stoer, 1966) to account for the mutual gravitational interaction between the $N=2$ black holes to 3.5PN order (Mora and Will, 2004). The integration is terminated if the BBHs get as close as twelve combined Schwarzschild radii. The BBHs are initialised on a highly eccentric orbit with eccentricity $e=0.999$, total mass $m=30\,\rm M_{\odot}$, and periapsis $r_{p}=300M$ where $M=Gm/c^{2}$. The mass ratio of the BBHs is drawn from a power-law distribution $p(q)\sim q^{\alpha}$ between 0.1 and 1.0 and the spin magnitudes from some Beta distribution $p(\chi_{1,2})\sim\rm Beta(a,b)$ with our fiducial parameters $\alpha=1.1$, $a=1.4$, and $b=3.6$. The BBHs are initialised at apoapsis and their spins are assumed to be aligned. Figure 6 shows that for the majority of BBH inspirals $\chi_{\rm eff}$ does not oscillate by more than $\sim 0.02$ and none of them by more than $\sim 0.1$ around the initial values. Based on these simulations, we conclude that our results are not significantly affected by treating $\chi_{\rm eff}$ as a conserved quantity.
@@ Appendix B Evolution of $\cos\theta_{1}$ and $\cos\theta_{2}$ during eccentric BBH inspirals
Unlike the effective spin parameter $\chi_{\rm eff}$, the spin tilt angles do not remain constant during the inspiral. Instead, they precess within precession cones whose amplitudes evolve due to GW emission, while remaining effectively fixed on the shorter precession timescale. The presence of distinct timescales in the dynamics allows for an efficient treatment of spin evolution through a double-averaging procedure: first over the orbital timescale and then over the precession timescale (Gerosa et al., 2015, 2023). This technique has been successfully applied to study spin-precessing, quasi-circular BBHs and, more recently, eccentric systems as well (Fumagalli and Gerosa, 2023). Using the extended version of the precession code presented by Fumagalli and Gerosa (2023), we investigate the evolution of the spin tilts in a population of eccentric BBHs in the mutually aligned and uncorrelated models as described in Sec. I. We simulate $10^{5}$ binaries with total mass $M=20M_{\odot}$; the mass ratios are drawn from a power-law distribution $p(q)\sim q^{\alpha}$ with $\alpha=1.1$ and $q\in[0.1,1.0]$; spin magnitudes are either sampled from a Beta distribution $p(\chi_{1,2})\sim\text{Beta}(a,b)$ with $a=1.4$ and $b=3.6$ or fixed to the maximal value $\chi_{1}=\chi_{2}=1$. BBHs are assumed to form with the initial eccentricity $e=0.999$ and semi-major axis $a\sim 1.1\times 10^{5}GM/c^{2}$, corresponding a GW frequency $f_{\rm GW}\sim 0.95$ Hz. We evolve the systems until they reach an eccentricity of $e=0.1$ at $f_{\rm GW}=10$ Hz.
Figure 7: Distribution of the cosine of spin tilt angles, $\cos\theta_{1}$ (left panels) and $\cos\theta_{2}$ (right panels). We assume spin tilts to be either mutually aligned (maroon) or uncorrelated (teal) at formation. Distributions at detection ($f_{\rm GW}=10$ Hz) and BBH formation ($f_{\rm GW}=0.95$ Hz) are shown as filled and dashed histogram, respectively. We assume either maximally spinning BH (top panels) or spin magnitudes extracted from a Beta distribution as in Fig. 2 (bottom panels).
Figure 7 shows the resulting evolution of the spin tilts $\theta_{1,2}$. As previously reported in quasi-circular binaries (Bogdanović et al., 2007; Gerosa et al., 2015), initially isotropic (i.e., uncorrelated) spin tilt distributions remain isotropic throughout the inspiral, even in the presence of eccentricity (Fumagalli and Gerosa, 2023). On the other hand, deviations appear in the case of initially aligned spin tilts. In these systems, the spin tilt distributions evolve away from mutual alignment, exhibiting a tendency to cluster around $\cos\theta_{i}=\pm 1$. This behavior, already identified in the quasi-circular case (Gerosa et al., 2013), is here found to persist in the eccentric regime as well. The effect is especially pronounced in the maximally spinning case, consistent with the expectation that spin-orbit and spin-spin couplings are stronger for higher spin magnitudes.